The following information was taken from the financial statements of the Frisco Corporation. All dollar amounts are in billions.

Assets, 12/31/00

$400

Common Stock, 12/31/00

$70

Dividends, 1/1/00-12/31/00

$8

Expenses, 1/1/00-12/31/00

$90

Liabilities, 12/31/00

$250

Retained Earnings, 12/31/00

$80

Revenues 1/1/00-12/31/00

$120

Determine the net dollar amount by which the Corporation's resources changed through management operations in 2000.

a.

$120 increase

b.

$80 increase

c.

$22 increase

d.

$30 increase
Answer #1

c.$22 increase

Revenues 1/1/00-12/31/00-Expenses, 1/1/00-12/31/00-Dividends, 1/1/00-12/31/00=120-90-8=$22